OpenStudy (anonymous):

Find the standard form of the equation of the parabola given vertex and two points.

OpenStudy (anonymous):

Vertex = (3,1) Point A = (2,0) Point B = (4,0)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Same approach should work

OpenStudy (anonymous):

How do I do it with 2 points? :/

OpenStudy (whpalmer4):

Ignore 1 of them? :-) No, they are a clue as to which way the parabola opens

OpenStudy (anonymous):

y = a(x-h)^2+k 0 = a(2-3)^2+1 0 = a+1 a = -1 y= -1(x-3)^2+1 I'm confused on what to do now :/

OpenStudy (whpalmer4):

Once you know that, you know which form to use, and you can use either point to fit the parabola

OpenStudy (whpalmer4):

Put in standard form and you are done

OpenStudy (anonymous):

y-1 = -1(x-3)^2 -y+1 = (x-3)^2 -(y-1) = (x-3)^2 would this be correct?

OpenStudy (whpalmer4):

Expand it to ax^2+bx+c Gotta go, good luck!

OpenStudy (anonymous):

thanks :D